### v2

# Near-real-time flood inundation monitoring by Bayesian analysis for change point problems for Sentinel-1 time series

Abstract
Near real-time flood monitoring is crucial for disaster response, yet existing methods face significant limitations in training data requirements and cloud cover interference. Here we present a novel approach using Bayesian analysis for change point problems (BCP) applied to Sentinel-1 SAR time series data, which automatically detects temporal discontinuities in backscatter patterns to distinguish flood inundation from permanent water bodies without requiring training data or ancillary information. We validate our method using the UrbanSARFloods benchmark dataset across three diverse geographical contexts (Weihui, China; Jubba, Somalia; and NovaKakhovka, Ukraine). Our BCP approach achieves F1 scores ranging from 0.41 to 0.76 (IoU: 0.25-0.61), significantly outperforming both OTSU thresholding (F1: 0.03-0.12, IoU: 0.02-0.08) and Siamese convolutional neural network approaches (F1: 0.08-0.34, IoU: 0.05-0.24). Further analysis reveals exceptional performance in open areas with F1 scores of 0.47-0.81 (IoU: 0.31-0.68) and high recall (0.36-0.84), contrasted with substantially lower performance in urban areas (F1: 0.00-0.01, IoU: 0.00-0.01), indicating a common challenge across current flood detection methods in urban environments. The proposed method’s ability to process raw SAR data directly with minimal preprocessing enables integration into operational early warning systems for rapid flood mapping, particularly in agricultural and open landscapes where it demonstrates the strongest performance.
keywords:
Synthetic Aperture Radar , Flood monitoring , Earth observation
journal: Remote Sensing of Environment
[…]
[not,tt,ns]organization=Graduate school of Science & Engineering, Saitama University, addressline=255 Shimo-Okubo, city=Sakura ward, Saitama, postcode=338-8570, state=Saitama, country=Japan
{graphicalabstract}[Uncaptioned image]
{highlights}
Novel Bayesian analysis for change point problems detects floods in Sentinel-1 time series
Method requires no training data or ancillary info to distinguish flood from permanent water
Automated processing of raw SAR data enables near real-time flood monitoring applications
1 Introduction
Flooding is one of the most frequent and devastating natural disasters globally [Alfieri2017-ma, Najibi2018-ta], resulting in substantial loss of life, damage to infrastructure, and severe economic consequences [Kundzewicz2014-ge, Jongman2021-ea, Tellman2021-hv, Jonkman2024-hi, Rentschler2021-dc]. The economic impact and human toll of flooding necessitate effective monitoring systems for early detection and rapid response [Oddo2019-or, Rentschler2020-pm]. Near real-time flood inundation mapping, enabled by advances in remote sensing, has emerged as a critical tool for effective disaster response [Oddo2019-or, Katiyar2021-fy, Shen2019-aj, Tay2020-cp]. These systems provide situational awareness to disaster responders, guide evacuation planning and search and rescue operations, and enable rapid damage assessments [Browder2021-io]. However, the widespread adoption of near real-time flood mapping remains limited due to challenges in remote sensing data analytics, creating an urgent need for more efficient and accessible monitoring approaches.
Satellite remote sensing offers significant advantages for flood monitoring, including wide spatial coverage, high temporal resolution, and the ability to observe remote or inaccessible areas. However, optical remote sensing techniques face substantial limitations due to their inability to penetrate cloud cover, which is often present during flood events. Synthetic Aperture Radar (SAR) addresses these limitations by operating in the microwave portion of the electromagnetic spectrum, enabling all-weather, day-night imaging capability. The launch of the Sentinel-1 mission by the European Space Agency (ESA) has greatly enhanced the availability of SAR data, providing global coverage with a revisit time of 6-12 days and free data access. With high spatial resolution (5-20 m) and wide swath width (250 km), Sentinel-1 is well-suited for mapping flood extents at regional to continental scales [Schlaffer2015-uv].
Despite SAR’s advantages for flood monitoring, significant challenges persist in developing fully automated, efficient, and accurate flood detection methods. Many existing approaches rely on supervised classification techniques that require extensive training data [Shastry2023-wl, Wu2023-kc, Ghosh2024-ub], which may not be readily available for all flood scenarios. Supervised classification methods require large amounts of labeled data to train algorithms, creating significant barriers to operational flood monitoring. The transferability of trained models to new flood scenarios or geographical regions is often limited [Zhao2024-lv], requiring additional data collection and model retraining. Furthermore, reliance on ancillary data (land cover maps, digital elevation models, hydrological information) can restrict the applicability and timeliness of flood mapping approaches, particularly in rapidly changing environments or regions with limited data infrastructure. Most existing time-series approaches for SAR-based flood monitoring rely either on simple change detection between pre- and post-flood images or on conventional statistical measures that don’t fully leverage the temporal information content in SAR data. Traditional change detection methods struggle to account for seasonal variations in backscatter or to distinguish between different types of change (e.g., flooding versus agricultural harvesting). While some Bayesian approaches have been applied in other remote sensing applications [Barry1993-qp, Erdman2008-kv], they have not been effectively adapted to exploit the unique characteristics of SAR time series for flood detection, particularly the need to identify abrupt changes against background variability.
Thus, this study aims to overcome these limitations by developing an efficient, automated solution for near real-time flood monitoring using Sentinel-1 SAR time series data. We propose a pixel-wise Bayesian statistical model for SAR-based flood mapping that operates without the need for training data. Unlike existing approaches, our method leverages a Bayesian analysis of Change Point problem (BCP) as proposed by Barry and Hartigan [Barry1993-qp] to identify flood-induced changes in the mean of our SAR time series data for the first time. Our workflow distinguishes between permanent water bodies and temporary flood inundations without requiring ancillary data by leveraging the temporal change characteristics in the probability domain rather than the raw backscatter domain. We use raw Sentinel-1 Ground Range Detected (GRD) time series as input without requiring training data, pre-trained models, or extensive manual intervention. The method achieves robust performance across diverse geographical contexts, making it particularly suitable for operational flood monitoring applications.
2 Related works
2.1 Flood monitoring by Optical remote sensing data
Optical satellite platforms such as Sentinel-2 and PlanetScope offer high spatial resolution and multispectral capabilities for flood mapping [Portales-Julia2023-pz, Wieland2019-je]. However, cloud cover significantly impedes continuous monitoring during critical initial flood phases. Portalés-Julià et al. (2023) found approximately 50% of pixels obscured by clouds during the first post-flood satellite revisit [Portales-Julia2023-pz]. Recent deep learning approaches have attempted to estimate water bodies beneath clouds [Li2021-fq, Shastry2023-wl]. These models leverage spatial context from adjacent cloud-free pixels and temporal dependencies between acquisitions, achieving over 90% accuracy in water classification under 50-80% cloud coverage [Portales-Julia2023-pz]. Despite these advances, challenges remain regarding training data requirements and geographical transferability.
2.2 Flood monitoring by remote sensing SAR data
SAR presents a compelling alternative due to its all-weather imaging capability [Landuyt2019-cz, Amitrano2024-av, Shen2019-aj]. SAR discriminates between water and non-water surfaces based on distinctive backscattering properties, with smooth water surfaces typically appearing dark due to specular reflection.
Interferometric SAR (InSAR) utilizes phase coherence information between acquisitions to detect inundation patterns, demonstrating effectiveness in complex urban environments [Chini2019-me]. However, the practical application of InSAR for timely flood monitoring is hindered by significant challenges. Normally, the extensive processing needed to derive these parameters often prevents real-time flood monitoring, limiting the method’s operational utility in urgent disaster response scenarios.
Deep learning methods have demonstrated exceptional capabilities for SAR-based flood detection [Saleh2024-vd, Bereczky2022-qy], with recent studies reporting classification accuracies exceeding 98% in certain contexts [Ghosh2024-ub]. Specialized datasets like Sen1Flood11 [Bonafilia2020-sk] and Kuro-Siwo [Bountos2023-jg] have facilitated model training across diverse geographical contexts.
Nevertheless, significant challenges persist: (1) extensive labeled training data requirements, (2) limited model transferability across geographical contexts [Wu2023-kc], (3) difficulty distinguishing between permanent water bodies and temporary inundation [Notti2018-vs], and (4) complex backscatter interpretation in urban or vegetated areas [Amitrano2024-av].
2.3 Time Series Analysis for Flood Detection
Recent research has increasingly leveraged the temporal dimension of satellite data for improved flood detection [Schlaffer2015-uv]. Time series approaches exploit the sequential nature of acquisitions to identify anomalous patterns indicative of flood events, particularly effective with Sentinel-1’s consistent 6-12 day revisit cycle [Oddo2019-or]. Schlaffer et al. (2015) proposed a harmonic analysis approach using multi-temporal SAR data to detect floods based on deviations from expected seasonal patterns [Schlaffer2015-uv]. Similarly, Tsyganskaya et al. (2018) and Tsyganskaya et al. (2019) employed time series analysis for wetland change detection and flooded vegetation mapping using Sentinel-1 data [Tsyganskaya2018-ts, Tsyganskaya2019-nv]. These methods enable more robust discrimination between permanent water bodies and temporary inundation.
Conventional Bayesian methods offer advantages for hydrological applications by incorporating prior knowledge and quantifying uncertainty [Barry1993-qp, Erdman2008-kv]. D’Addabbo et al. (2016) developed a Bayesian network model fusing SAR and ancillary data [D-Addabbo2016-yi], and Giordano et al. (2023) applied Bayesian statistical modeling to multi-temporal Sentinel-1 stacks for high-resolution flood monitoring [Giordano2023-ys]. However, existing Bayesian approaches primarily focus on retrospective analysis rather than near real-time detection. Most methods lack specific adaptations for the characteristic SAR backscatter patterns during flood events—sharp decreases during inundation followed by gradual increases during recession, limiting their effectiveness for timely flood monitoring.
2.4 Research Gap and Our Contribution
Previous studies faces challenges in SAR-based flood monitoring that our approach addresses: (1) extensive training data requirements limiting applicability in data-scarce regions, (2) domain adaptation difficulties across flood scenarios, and (3) challenges distinguishing permanent water bodies from flood inundation. Our approach with Bayesian analysis for change point problems (BCP) operates without training data, processes Sentinel-1 time series directly, and leverages temporal patterns to differentiate water classes without pre-existing water masks. This approach fills a critical gap in flood monitoring studies, offering a complementary method particularly valuable when timely detection is essential for emergency response.
3 Methods
The workflow of our approach consists of three main components: (1) Reference data and Sentinel-1 time series data preparation, (2) BCP analysis, and (3) post-processing with spatial filtering. The workflow is designed to process raw SAR data directly and provide automated flood extent mapping without requiring training data or extensive manual intervention.
3.1 Data Preparation
3.1.1 UrbanSARFloods
As reference data, we utilized the UrbanSARFloods dataset, a comprehensive benchmark for flood detection using Sentinel-1 SAR data [Zhao2024-lv]. This dataset encompasses 8,879 non-overlapping image chips of 512×512 pixels at 20-m spatial resolution, covering an area of 807,500 km2 across 18 flood events on 5 continents. Every pixel in this dataset is classified into three distinct categories: non-flooded areas (0), flooded open areas (1), and flooded urban areas (2). UrbanSARFloods is systematically divided into training (4,501 chips), validation (1,970 chips), and testing (2,408 chips) sets, which are independent of flood events in training and validation data. This dataset ensures its quality by a two-tier annotation process, which combines semi-automatic labeling through conventional remote sensing techniques with manual verification using high-resolution optical data from PlanetScope and UAV imagery. While the BCP analysis does not require any training/validation set, we used them for the comparative analysis described in Section 4.
To evaluate the geographical robustness of our method, we conducted extensive testing across three diverse test sites found in UrbanSARFloods: Weihui (China), Jebba (Somalia), and NovaKakhovka (Ukraine). Weihui represents an urban environment with complex infrastructure and dense building patterns. In such settings, SAR backscatter is influenced by double-bounce reflection mechanisms between water surfaces and vertical structures, creating complex scattering patterns that can confound conventional flood detection approaches. The regular grid pattern of roads and buildings creates distinctive backscatter signatures that require careful interpretation. Jebba exemplifies a rural, agricultural landscape with varying topography and seasonal vegetation patterns. In these areas, the temporal variability of vegetation growth can mask or mimic flood signatures, particularly in partially submerged vegetation where volume scattering mechanisms dominate. The predominantly open terrain presents different challenges compared to urban environments, with larger homogeneous regions but greater seasonal variability. NovaKakhovka represents a mixed environment with both urban elements and natural landscapes, including significant riverine features. This combination of built environment and natural terrain presents a particularly challenging test case, as algorithms must correctly distinguish permanent water bodies (river) from newly flooded areas while accounting for different backscatter behaviors across varied land cover types.
3.1.2 Sentinel-1 GRD
Sentinel-1 GRD offers VV and VH polarization modes that measure radar backscatter amplitude using C-band SAR instruments. The data undergoes multi-looking processing and is projected onto the WGS84 ellipsoid, converting radar measurements into ground coordinates for geographic analysis. This multi-looking process reduces noise while preserving backscatter amplitude, though it removes phase information. For each flood event, we collected time series data spanning one year before the event using Google Earth Engine, to establish a robust baseline for change detection. We didn’t take into account of further processing such as speckle filtering and radiometric terrain normalization to ensure the sensitivity of time-series variability of backscatter intensity to changes in surface properties. To account for differences in viewing geometry and illumination conditions over time, we filtered the timeseries data in the platform type (A or B) and the orbit direction (ascending or descending) that influences backscatter response due to varying incidence angles and shadow effects based on the observation at the flood event (Table 1). While Sentinel-1 GRD data is 10-m spatial resolution, we aggregated into 20-m spatial resolution to match with UrbanSARFloods in this experiment.
Table 1: Tested flood events and observation condition during the flood event.
Selected event country Orbit Day of image for annotation
Weihui China Ascending 27th July, 2021
Jebba Somalia Decending 1st December, 2023
NovaKakhovka Ukraine Ascending 9th June, 2023
3.2 Bayesian analysis for change point problems
The BCP assumes the observations $X_{1},...,X_{n}$ are independent and normally distributed with $X_{i}\sim N(\mu_{i},\sigma^{2})$. The probability of a changepoint at position $i$ is denoted by $p$ and is assumed to be independent at each $i$. The prior distribution of the mean $\mu_{ij}$ for the block from position $i+1$ to $j$ is chosen to be $N(\mu_{0},\sigma_{0}^{2}/(j-i))$. This prior allows the detection of subtle changes given sufficient data.
The BCP uses a partition $\rho=(U_{1},...,U_{n})$, where $U_{i}=1$ indicates a changepoint at position $i+1$. At each step of the Markov chain Monte Carlo (MCMC) procedure, a value of $U_{i}$ is drawn from the conditional distribution of $U_{i}$ given the data and current partition. The transition probability $p$ for the conditional probability of a changepoint at position $i+1$ is calculated as:
[…]
where $b$ is the number of blocks if $U_{i}=0$, $W_{0}$ and $B_{0}$ are the within and between block sums of squares when $U_{i}=0$, $W_{1}$ and $B_{1}$ are the equivalent when $U_{i}=1$, and $\gamma$ and $\lambda$ are tuning parameters.
The within block sum of squares $W$ is calculated as:
$W=\sum_{ij\in\rho}\sum_{l=i+1}^{j}(X_{l}-\bar{X}_{ij})^{2}$ (2)
where $\bar{X}_{ij}$ is the mean of the observations in the block from $i+1$ to $j$.
The between block sum of squares $B$ is calculated as:
$B=\sum_{ij\in\rho}(j-i)(\bar{X}_{ij}-\bar{X})^{2}$ (3)
where $\bar{X}$ is the overall mean of the data.
We implemented BCP using the bcp R package [Erdman2008-kv]. We selected the default tuning parameters of $\gamma=\lambda=0.2$ as recommended by Barry and Hartigan [Barry1993-qp, Erdman2008-kv]. Posterior means and probabilities were calculated based on 500 MCMC iterations after discarding the first 50 as burn-in, a configuration that achieved convergence while maintaining computational efficiency.
To our SAR time series data over a 1-year period prior to each flood event, we use the posterior probability of a changepoint at the last observation during the flood as input for the subsequent processing. This is expected to detect potential changes in the SAR signal associated with the flood while accounting for normal seasonal variations in backscatter. Three different sets of inputs to the BCP analysis were considered to evaluate the sensitivity and performance of our approach: (1) VV polarization only, (2) VH polarization only, and (3) VV-VH combined. The VV polarization is generally more sensitive to surface roughness and typically provides stronger contrast between water and non-water surfaces. The VH cross-polarization is often more sensitive to volume scattering and can provide complementary information, particularly in vegetated areas. The combined VV-VH approach leverages information from both polarizations, potentially enhancing the detection capabilities across diverse environmental contexts. By examining these three input configurations, we assessed which combination of polarimetric information most effectively captured flood-induced changes in the SAR backscatter time series.
3.3 De-noising
Sentinel-1 SAR images inherently contain multiplicative speckle noise which may lead to false detections of flooded areas. To reduce this noise while preserving spatial details, we applied a local median filter to the results of the BCP analysis. A local median filter with a window size of $w$ x $w$ pixels was applied to the BCP output. The window size $w$ is a parameter that can be adjusted based on the spatial resolution of the input imagery and the desired level of smoothing. After applying the filter, the smoothed values were thresholded using a parameter $t$. Pixels with values above this threshold were considered as potential flood areas, while those below were masked out.
To determine optimal parameters for our flood detection method, we conducted a systematic parameter sensitivity analysis focusing on the $w$ and $t$. For the window size parameter, we explored values ranging from 3 to 15 pixels with a step size of 2. The lower bound of 3 represents the minimum window size needed for effective noise reduction while the upper bound of 15 was chosen to avoid excessive smoothing that could eliminate small but significant flooded areas. The 20 m spatial resolution of our analysis data meant that a window size of 3 would cover an area of 60m × 60m on the ground, while a window size of 15 would cover 300 m × 300 m.
For the threshold parameter ($t$), we investigated values between 0.1 and 0.9 with a step size of 0.1. This comprehensive range covers the entire probability space from very permissive detection (0.1) to highly selective identification (0.9). Theoretically, changepoint probabilities above 0.5 indicate a more likely than not change, but in practice, lower thresholds may be effective for detecting subtle hydrological changes. Thus, the parameter optimization process involved evaluating all 63 combinations of window sizes and thresholds (7 window sizes × 9 threshold values) for each flood event (local) or across multiple flood events (global) in diverse geographical settings to ensure the robustness and transferability of our approach.
4 Evaluation
To demonstrate the effectiveness of our proposed approach, we compared our results with both conventional and modern deep-learning approaches for flood detection.
4.0.1 OTSU Thresholding Method
As a conventional baseline, we implemented the Otsu thresholding method [Otsu1979-dj]. This non-parametric technique determines an optimal threshold by maximizing the between-class variance between flood and non-flood pixels in SAR imagery. We applied the Otsu method to two different input configurations: VV polarization only and VH polarization only.
The Otsu algorithm operates by computing the weighted sum of variances of the two classes:
$\sigma_{w}^{2}(t)=\omega_{0}(t)\sigma_{0}^{2}(t)+\omega_{1}(t)\sigma_{1}^{2}(t)$ (4)
where $\omega_{0}$ and $\omega_{1}$ represent the class probabilities, $\sigma_{0}^{2}$ and $\sigma_{1}^{2}$ denote the class variances, and $t$ is the threshold value. We followed standard practices in remote sensing literature for the Otsu method, including histogram equalization preprocessing to enhance contrast in the SAR imagery before applying the threshold calculation.
4.0.2 Siamese CNN Method
As a representative of modern deep learning approaches, we built a Siamese Convolutional Neural Network (CNN) from the training and validation sets of the UrbanSARFloods dataset. The network architecture consists of dual processing branches with shared weights to compare the latest observation with the second latest observation. The model takes two input images of size 512×512 pixels with 2 channels each (VV and VH polarizations): Input A (latest observation) and Input B (second latest observation). Each image is processed through identical convolutional branches with shared weights, ensuring that features are extracted using the same transformations. The encoder pathway implements sequential convolutional blocks with increasing filter depths (32, 64, and 128 filters), where each block contains:
  • 1.
    3×3 convolution with ’same’ padding to maintain spatial dimensions
  • 2.
    Batch normalization to stabilize and accelerate training
  • 3.
    ReLU activation function to introduce non-linearity
The decoder pathway incorporates skip connections from corresponding encoder layers to preserve spatial details through addition operations rather than concatenation. A custom subtraction layer computes the difference between the processed features from both branches to identify changed regions, followed by a 1×1 convolution with sigmoid activation to produce the final flood probability map.
The model was trained using binary cross-entropy loss function, with consideration for class imbalance where the positive class weight was calculated to be approximately 35.67. We used the Adam optimizer with an initial learning rate of 0.001 and gradient clipping (clipnorm=1.0), along with a ReduceLROnPlateau scheduler that reduced the learning rate by half after 5 epochs of no improvement. The model was trained for up to 200 epochs with early stopping (patience=20) based on validation loss.
The training dataset comprised 231 image pairs with 58 pairs reserved for validation from urban flooding events. Each input was normalized using Z-score normalization and clipped to ±3 standard deviations to handle outliers. NaN values were replaced with zeros to ensure numerical stability.
4.0.3 Accuracy assessment
We evaluated our method’s performance using standard metrics including precision, recall, F1 score, and Intersection over Union (IoU). The evaluation was conducted in two ways: first as an overall flood detection assessment, and then separately for urban and open areas to account for different challenges in each context. For the overall flood detection evaluation, we considered both flood open and urban pixels as positive cases. For the separate analysis, we evaluated the performance independently for each land cover type according to the reference data.
Precision represents the proportion of correctly predicted flood water pixels among all positive predictions, while recall measures the proportion of actual flood water pixels that were correctly identified. Precision and recall are calculated as follows:
$Precision=\frac{N_{TP}}{N_{TP}+N_{FP}}$ (5)
$Recall=\frac{N_{TP}}{N_{TP}+N_{FN}},$ (6)
where $N_{TP}$ represents the number of true positives (correctly predicted flood water pixels), $N_{FP}$ represents the number of false positives (incorrectly predicted as flood water), and $N_{FN}$ represents the number of false negatives (flood water pixels not detected).
The F1 score, which provides a balanced measure of the model’s performance, is calculated as:
$F1=2\times\frac{Precision\times Recall}{Precision+Recall}$ (7)
The Intersection-over-Union (IoU) was aimed to quantify the spatial overlap between the predicted flood water regions and the ground truth annotations:
$IoU=\frac{Area\ of\ Overlap}{Area\ of\ Union}=\frac{N_{TP}}{N_{TP}+N_{FP}+N_{FN}}$ (8)
For both the overall and separate evaluations, these metrics were calculated for each test site (Weihui, Jubba, and NovaKakhovka).
5 Results
5.1 BCP Outputs
The BCP analysis outputs revealed that the estimated flood extents are sensitive to the input variables (Figure 1). The visual comparison of the BCP outputs suggests that the BCP results using VH alone seem less capable of detecting flood inundation than the results using VV or VV-VH input. The use of VV-VH input delineates the flood extent more clearly than the others at all three sites.
Figure 1: Randomly selected patches for: Jubba (Somalia) (a) Bing map that represent non-flood condition, (b) Referenced UrbanSARFloods data set, (c) BCP results used VV input only, (d) BCP results used VH input only, (e) BCP results used VV-VH input; Weihui (China) (f) Bing map that represent non-flood condition, (g) Binary UrbanSARFloods data set, (h) BCP results used VV input only, (i) BCP results used VH input only, (j) BCP results used VV-VH input; NovaKakhovka (Ukraine) (k) Bing map that represent non-flood condition, (l) Binary UrbanSARFloods data set, (m) BCP results used VV input only, (n) BCP results used VH input only, (o) BCP results used VV-VH input.
5.2 Parameter Sensitivity Analysis
The original BCP results are the posterior probabilities of flood occurrence at each pixel. To delineate the flood extents, we post-processed the BCP results with different parameter settings for threshold and window size and created the binary (flood/non-flood) outputs. Considering the result using VV-VH input only, as this is thought to be the best performance for the BCP, Figure 2 shows the comparison of the binary outputs with different parameter settings for Jubba (Somalia) as an example. The larger $w$ tends to overestimate the flood extent, while the smaller $w$ may fail to omit the speckle noise. The larger $t$ picks up only the flood extent with the higher confidence, while the smaller $t$ may include false positives.
Figure 2: Comparative maps of a randomly selected patch for Jubba (Somalia) with different parameter settings for post-processing.
To evaluate the performance of the estimated flood extents, we explored the parameter sensitivity based on the F-1 score against threshold and window (Figure 2). The optimal parameter combination emerged as a threshold value of 0.2 and a window size of 9 pixels, achieving a mean F1-score of 0.60 across all test sites. However, performance varied substantially across locations, with the highest performance at Weihui (F1-score: 0.76), followed by moderate performance at Jubba (0.65), and lower performance at NovaKakhovka (0.41).
Figure 3: Parameter sensitivity analysis of flood detection method based on F-1 score against threshold and window size in Weihui (China), Jubba (Somalia), and NovaKakhovka (Ukraine).
We identified 44 of 63 parameter combinations that showed no statistically significant difference from the optimal configuration ($p>0.05$). These combinations formed several key clusters: combinations with $t=0.2$ and window sizes 7, 11, and 13 (Mean F1 $>$ 0.59, see table 2); combinations with $t=0.3$ and window sizes 5-15 (Mean F1 $>$ 0.56); and some combinations with $t=0.1$ (Mean F1 $>$ 0.50).
Performance degradation became statistically significant for threshold values $\geq$ 0.4, particularly when combined with extreme window sizes (Figure 3). The decline in performance showed a consistent pattern with increasing threshold values and window size.
Table 2: Top performing parameter combinations and their statistical significance. Selected results only with t=0.2.
Threshold (t) Window Size (w) Mean F1-Score Std F1-Score p-value
0.2 9 0.60 0.18 -
0.2 7 0.60 0.18 0.69
0.2 11 0.60 0.18 0.22
0.2 13 0.60 0.18 0.06
0.2 5 0.59 0.16 0.43
5.3 Comparative Analysis Across Sites
With the use of the optimal parameters for BCP, Figure 4 illustrates the comparison of our method with the conventional and deep-learning approaches across randomly selected three example patches. Each column represents a different study area, with rows showing the referenced UrbanSARFloods data, OTSU thresholding (VV, VH), Siamese CNN, and BCP (VV, VH, VV-VH) with the optimized parameters. The OTSU thresholding approach seems a good performance to Wuhei, but it tends to overestimate the flood extent in Jubba and NovaKakhovka. VV input seems to perform better than VH input in all three sites in terms of noise reduction. Siamese CNN shows a good performance and successfully overcomes the noise compared to the other approaches. However, it tends to underestimate the flood extent in Jubba and NovaKakhovka. BCP with VV or VH alone tends to have noise in the flood extent, while the BCP with VV-VH input shows relatively good in terms of noise reduction.
Figure 4: Flood detection results across randomly selected patches in different geographical contexts and estimated approaches.
5.4 Quantitative Performance Evaluation
The table LABEL:tab:comparison summarizes the accuracy measures of our approach compared to OTSU thresholding and Siamese CNN approaches. For this evaluation, we used the globally optimazed parameters identified through sensitivity analysis (t=0.2, w=9) with VV-VH input. Our BCP approach achieved F1 scores ranging from 0.41 to 0.76, significantly outperforming OTSU thresholding and Siamese CNN approaches across all sites. The OTSU thresholding approach struggled with false positives, resulting in low precision across all sites. While the Siamese CNN showed improved performance over OTSU, it still fell short of our method’s capabilities.
Table 3: Accuracy assessment of flood detection methods across three flood events.
Weihui Jubba NovaKakhovka
Precision Recall F1 IoU Precision Recall F1 IoU Precision Recall F1 IoU
OTSU (VV) 0.02 0.37 0.03 0.02 0.08 0.49 0.11 0.07 0.03 0.58 0.05 0.03
OTSU (VH) 0.03 0.43 0.04 0.03 0.09 0.57 0.12 0.08 0.03 0.54 0.05 0.03
Siamese-CNN 0.33 0.42 0.34 0.24 0.38 0.26 0.22 0.16 0.18 0.11 0.08 0.05
BCP 0.73 0.79 0.76 0.61 0.57 0.74 0.64 0.48 0.63 0.30 0.41 0.25
5.5 Analysis of Performance Variations Between Urban and Open Areas
To better understand the performance characteristics across different land cover types, we conducted separate analyses for flood urban and flood open pixels. Tables 4 and 5 present the detailed accuracy metrics for each category and method. In open areas, all methods demonstrated measurable effectiveness in flood detection, though with varying degrees of success. The BCP method achieved the highest performance in Jubba with a precision of 0.37 and recall of 0.75 (IoU: 0.37), followed by strong results in Weihui (precision: 0.31, recall: 0.62, IoU: 0.31). However, its performance declined significantly in NovaKakhovka (precision: 0.05, recall: 0.09, IoU: 0.05). The OTSU method showed consistent but modest performance across sites, with recall values ranging from 0.21 to 0.56, though precision remained low (0.01-0.08). The Siamese CNN demonstrated more balanced performance in open areas, particularly in Weihui where it achieved an IoU of 0.13, though its effectiveness varied considerably across sites.
The analysis of urban environments revealed more challenging conditions for flood detection across all methods. The conventional OTSU method exhibited extremely low precision (0.000-0.002) despite achieving moderate recall values (0.02-0.23), resulting in poor overall performance with F1 scores below 0.003. The Siamese CNN demonstrated notably superior performance in urban areas compared to other methods, achieving precision values ranging from 0.11 to 0.32 and recall values between 0.12 and 0.17. This translated to F1 scores between 0.12 and 0.16, with the highest performance observed in Jubba (IoU: 0.10) and NovaKakhovka (IoU: 0.11). The BCP method, while excelling in open areas, showed limited effectiveness in urban environments with precision below 0.005 and recall under 0.06 across all sites.
Table 4: Accuracy assessment of flood detection methods across three flood events for open areas.
Weihui Jubba NovaKakhovka
Precision Recall F1 IoU Precision Recall F1 IoU Precision Recall F1 IoU
OTSU (VV) 0.03 0.45 0.05 0.03 0.09 0.52 0.15 0.08 0.04 0.63 0.07 0.04
OTSU (VH) 0.04 0.51 0.07 0.04 0.10 0.61 0.17 0.09 0.04 0.59 0.07 0.04
Siamese-CNN 0.38 0.53 0.44 0.28 0.42 0.32 0.36 0.22 0.22 0.13 0.16 0.09
BCP 0.78 0.84 0.81 0.68 0.63 0.80 0.70 0.54 0.68 0.36 0.47 0.31
Table 5: Accuracy assessment of flood detection methods across three flood events for urban areas.
Weihui Jubba NovaKakhovka
Precision Recall F1 IoU Precision Recall F1 IoU Precision Recall F1 IoU
OTSU (VV) 0.00 0.23 0.00 0.00 0.00 0.09 0.00 0.00 0.00 0.02 0.00 0.00
OTSU (VH) 0.00 0.22 0.00 0.00 0.00 0.09 0.00 0.00 0.00 0.02 0.00 0.00
Siamese-CNN 0.11 0.17 0.12 0.08 0.32 0.12 0.14 0.10 0.26 0.17 0.16 0.11
BCP 0.01 0.06 0.01 0.01 0.00 0.03 0.00 0.00 0.00 0.02 0.00 0.00
6 Discussion
6.1 Methodological Advances in SAR-Based Flood Detection
Our study introduces Bayesian analysis for change point problems (BCP) to near real-time flood monitoring applied to Sentinel-1 GRD time series data. Unlike supervised methods that require extensive training data and face challenges with geographical transferability [Wu2023-kc, Ghosh2024-ub], our approach leverages the statistical properties of SAR time series to detect flood-induced changes without any training requirements. This fundamental training-free characteristic is a significant advantage, as it completely eliminates the overfitting issues inherent in deep learning and other machine learning approaches. Machine learning methods often perform well in areas represented in training datasets but fail when applied to novel geographical contexts or unusual flood patterns. By contrast, our BCP method requires no supervised learning process, making it immediately applicable to any region with sufficient SAR time series data, regardless of whether similar flood events have been previously observed and labeled.
Our method advances the field by reframing flood detection as a temporal changepoint identification problem rather than a binary classification task. Conventional classification approaches typically rely on absolute backscatter values or simple change metrics between two images, making them vulnerable to variations in instrument calibration, incidence angle effects, and seasonal changes in land cover [Schlaffer2015-uv]. In contrast, our Bayesian framework models the entire temporal trajectory of backscatter values, enabling more robust discrimination between normal fluctuations and flood-induced changes.
Parameter sensitivity analysis showed that a window size of 9 pixels provided a good balance between noise reduction and preservation of flood boundary details, though our method demonstrated relatively stable performance across a range of parameter settings.
6.2 Performance in Relation to State-of-the-Art Approaches
Our performance analysis reveals distinct patterns across different environmental contexts and detection methods. In open areas, our BCP method demonstrated robust performance with F1 scores ranging from 0.74 to 0.84, outperforming the OTSU thresholding method (F1: 0.031-0.132). The Siamese CNN showed variable performance in open areas, achieving its best results in Weihui (F1: 0.626) but performing less effectively in Jubba (F1: 0.366) and NovaKakhovka (F1: 0.173). This variability suggests that while deep learning approaches can achieve moderate success in open areas, they may be sensitive to regional differences in landscape characteristics and flood patterns.
The stark performance disparity between open and urban areas can be attributed to fundamental differences in SAR backscatter characteristics. Open areas offer more uniform backscatter patterns in SAR imagery, enabling more reliable flood detection across all methods due to clearer contrast between flooded and non-flooded surfaces. In contrast, urban environments present complex scattering mechanisms due to built infrastructure, leading to increased signal ambiguity. The presence of various materials and structures in urban areas creates multiple reflection pathways, making flood detection particularly challenging. These physical constraints fundamentally limit the effectiveness of current detection methods, regardless of their computational sophistication.
6.3 Integration with Operational Flood Monitoring Systems
The practical application of our method extends beyond academic research to operational flood monitoring systems. Our approach could complement existing near real-time flood mapping services such as the Copernicus Emergency Management Service (EMS) and NASA’s MODIS Near Real-Time Global Flood Mapping system. These systems currently rely primarily on optical imagery or simple thresholding of SAR data, facing challenges with cloud cover and discrimination between permanent and temporary water features.
For regional flood monitoring centers with limited computational resources, our method offers particular advantages. Unlike deep learning approaches that typically require GPU acceleration and substantial training data, our statistical framework operates efficiently on standard computing hardware and requires no training phase. This accessibility could enable more widespread adoption of SAR-based flood monitoring in developing regions where flood impacts are often most severe but technical resources are limited.
The method’s ability to process raw Sentinel-1 GRD data directly also streamlines the operational workflow, reducing the need for extensive preprocessing chains that can introduce delays in time-sensitive emergency response scenarios. Combined with Sentinel-1’s systematic acquisition plan, this could enable near-real-time flood alerts within hours of satellite overpass, a critical improvement over current systems that often require days to deliver actionable information.
6.4 Limitations and Future Research Directions
Despite the promising results, several limitations of our approach warrant consideration. The method’s performance in open areas, particularly in NovaKakhovka, indicates challenges in distinguishing flood inundation from other changes in agricultural regions. This limitation appears related to the complex temporal patterns of vegetation growth and harvest activities that can create backscatter changes similar to flooding. Future research could address this by incorporating ancillary information on land cover and seasonal agricultural calendars to better contextualize detected changes.
From a methodological perspective, the current BCP implementation assumes normally distributed backscatter values, which may not fully capture the complex statistical properties of SAR data. Alternative Bayesian frameworks that incorporate more sophisticated distributions, such as the gamma distribution commonly used for SAR intensity modeling, could potentially improve detection accuracy. Similarly, hierarchical Bayesian models that explicitly account for spatial context could enhance the method’s robustness to speckle noise without requiring post-processing spatial filtering.
The temporal resolution of Sentinel-1 (6-12 days) represents another limitation for monitoring rapidly evolving flood events [Munasinghe2023-oc]. Future research could explore the integration of our approach with other SAR platforms (e.g., RADARSAT Constellation Mission, ALOS-2) to achieve higher temporal frequency. Alternatively, data fusion techniques that combine SAR with optical or passive microwave observations could provide more continuous monitoring capabilities, leveraging the complementary strengths of different sensor types.
7 Conclusions
We present a novel approach to near real-time flood monitoring using Bayesian Changepoint Detection (BCP) with Sentinel-1 SAR time series data. Our key contributions include: (1) a training-free framework for automated flood detection, (2) innovative application of Bayesian analysis to SAR time series, and (3) empirical validation across diverse contexts.
As climate change increases flood risks globally, our method contributes to disaster resilience and adaptation efforts through its accessibility, transferability, and operational efficiency.
Our approach can be incorporated into semi-real-time flood monitoring systems due to its automatic processing capabilities. However, a key constraint is the infrequent coverage of Sentinel-1, which creates a bottleneck during urgent disaster response. Further research is needed to demonstrate our approach with higher-frequency SAR time series data.
8 Acknowledgements
This work was funded by NEDO (23200859-0).